# MSegRNN:Enhanced SegRNN Model with Mamba for Long-Term Time Series Forecasting

### v1

@@ 4.2 Implicit Segmentation
The explicit segmentation approach employed by SegRNN pre-processes the data such that a given multivariate time series, denoted as $X^{(i)}\in\mathbb{R}^{L\times C}$, where $L$ represents the look-back length and $C$ represents the number of channels in multivariate time series. Upon selecting a segment length $w$, the time series can be transformed into segmented time series $X_{w}^{(i)}\in\mathbb{R}^{n\times w\times C}$, where $n$ represents the number of segments:
@@
Notably, our proposed method does not explicitly define the segmentation length. Instead, it achieves data segmentation and dimensionality transformation through dual linear projections. The incorporation of linear fully connected layers facilitates a more continuous transformation of information. It is importanct to emphasize that this implicit segmentation strategy is not confined to a singe approach but rather represents a general and straightforward technique for information enhancement.
@@ 4.3 Residual Structure for SegRNN
Using segmentation techniques, SegRNN reduces the number of iterations from $L$ to $n$. However, the number of iterations still increases with the forecasting time steps. This unavoidably limits the model’s ability due to the loss of information during iterations. Therefore, we introduce a residual structure in the model. For the segmented vector $\bar{\mathbf{X}}^{(i)}\in\mathbb{R}^{C\times L\times n}$, we introduce a linear mapping from $\bar{\mathbf{X}}^{(i)}$ to $\mathbf{h}_{n}$. Specifically, we first merge the last two dimensions and extend the first dimension, obtaining the three-dimensional matrix $\mathbf{X}^{\prime(i)}\in\mathbb{R}^{1\times C\times(L\times n)}$. Then, we transform the last dimension through a linear layer to get $\mathbf{h}^{\prime}_{n}\in\mathbb{R}^{1\times C\times d}$. Finally, we add $\mathbf{h}_{n}$ to the corresponding position in $\mathbf{h}^{\prime}_{n}$, obtaining the final output vector from the encoder.
@@ 5 Experiments
Our experiments are structured as follows. In Section 5.1, we introduce the datasets used for long-term time series analysis. Subsequently, in Section 5.2, we provide a detailed description of the experimental setup and baseline models for comparison. Section 5.3 presents and analyzes the performance metrics. In Section 5.4, we conducted ablation studies to investigate the effectiveness of the model. Finally in Section 5.5, we investigated the efficiency of the model, particularly focusing on its memory consumption to demonstrate the lightweight characteristic of our model. All experiments in this section were conducted on a single NVIDIA 4090 GPU.
@@ 5.2 Experimental setup and Baselines
The unified configuration of the model is substantially aligned with the SegRNN approach. The look-back window is set to be 96, implicit Segmentation length is set to be 24 or 48 based on the size of datasets. A single layer of Mamba is utilized for preprocessing the data and a single GRU layer is used for sequence processing. The dimensionality of hidden layer with GPU structure is set to be 512, and the training epochs are set to be 30. Dropout rate, learning rate, and batch size vary with the data and the scale of the model.
@@ 5.3 Main Result
Our method achieved top positions in 37 out of 48 metrics, while the SegRNN method achieved the top position in 7 metrics. The iTransformer, Crossformer, and PatchTST methods each achieved top position in 6, 1 and 1 metrics, respectively. This demonstrates the powerful capabilities of our model.
@@
Count 37 7 6 1 1 0 0
@@ 5.4 Ablation study
720 0.341 0.340 0.350 0.342 0.380 0.377 0.357 0.376
Average 0.241 0.266 0.246 0.268 0.280 0.303 0.251 0.298
@@ 5.5 Model efficiency
Since our Mamba structure has not undergone hardware-level optimization, this section focuses on comparing the model’s memory usage to verify its lightweight characteristic. The lightweight characteristic is primarily due to the use of a small number of parameters in our Mamba structure experiments. Using the Weather dataset as an example, the memory usage comparison with batch size 64 is shown in Figure 2.
Figure 2: Mean memory consumpation in weather datasets for MSegRNN and SegRNN
As shown in the figure, the model’s exhibits only marginal increases in memory usage compared to SegRNN, affirming its lightweight characteristics.

### v2

@@ 4.2 Implicit Segmentation
The explicit segmentation approach employed by SegRNN pre-processes the data such that a given multivariate time series, denoted as $X^{(i)}\in\mathbb{R}^{L\times C}$, As shown in Figure 2, where $L$ represents the look-back length and $C$ represents the number of channels in multivariate time series. Upon selecting a segment length $w$, the time series can be transformed into segmented time series $X_{w}^{(i)}\in\mathbb{R}^{n\times w\times C}$, where $n$ represents the number of segments:
@@
Figure 2: Comparsion of SegRNN Segmentation with Implicit Segmentation
Notably, our proposed method does not explicitly define the segmentation length. Instead, it achieves data segmentation and dimensionality transformation through dual linear projections. The incorporation of linear fully connected layers facilitates a more continuous and dense transformation of information. It is importanct to emphasize that this implicit segmentation strategy is not confined to a singe approach but rather represents a general and straightforward technique for information enhancement.
@@ 4.3 Residual Structure for SegRNN
Using segmentation techniques, SegRNN reduces the number of iterations from $L$ to $n$. However, the number of iterations still increases with the forecasting time steps. This unavoidably limits the model’s ability due to the loss of information during iterations. Therefore, we introduce a residual structure in the model, as shown in Figure 3. For the segmented vector $\bar{\mathbf{X}}^{(i)}\in\mathbb{R}^{C\times L\times n}$, we introduce a linear mapping from $\bar{\mathbf{X}}^{(i)}$ to $\mathbf{h}_{n}$. Specifically, we first merge the last two dimensions and extend the first dimension, obtaining the three-dimensional matrix $\mathbf{X}^{\prime(i)}\in\mathbb{R}^{1\times C\times(L\times n)}$. Then, we transform the last dimension through a linear layer to get $\mathbf{h}^{\prime}_{n}\in\mathbb{R}^{1\times C\times d}$. Finally, we add $\mathbf{h}_{n}$ to the corresponding position in $\mathbf{h}^{\prime}_{n}$, obtaining the final output vector from the encoder.
Figure 3: Residual block for SegRNN
@@ 5 Experiments
Our experiments are structured as follows. In Section 5.1, we introduce the datasets used for long-term time series analysis. Subsequently, in Section 5.2, we provide a detailed description of the experimental setup and baseline models for comparison. Section 5.3 presents and analyzes the performance metrics. In Section 5.4, we conducted ablation studies to investigate the effectiveness of the model. Finally in Section 5.5, we investigated the efficiency of the model, particularly focusing on its memory consumption to demonstrate the lightweight characteristic of our model. All experiments in this section were conducted on two NVIDIA T4 GPU, The source code is available at https://github.com/UniqueoneZ/MSegRNN.git.
@@ 5.2 Experimental setup and Baselines
The unified configuration of the model is substantially aligned with the SegRNN approach. The look-back window is set to be 96, a single layer of Mamba is utilized for preprocessing the data and a single GRU layer is used for sequence processing. The dimensionality of hidden layer with GPU structure is set to be 512, and the training epochs are set to be 30. Dropout rate, learning rate, and batch size vary with the data and the scale of the model.
@@ 5.3 Main Result
Our method achieved top positions in 36 out of 48 metrics, while the SegRNN method achieved the top position in 7 metrics. The iTransformer, Crossformer, and PatchTST methods each achieved top position in 6, 1 and 1 metrics, respectively. This demonstrates the powerful capabilities of our model.
@@
Count 36 7 6 1 1 0 0
We subsequently conducted experiments to observe how the model’s capability varies with changes in the look-back window. The hyperparameter settings of the method were kept consistent with look-back window 96, and the horizon was uniformly set to 336. The evaluation was performed on the ETTh2 dataset, with the MSE and MAE losses of the method shown in the Figure 4:
Figure 4: Comparsion of SegRNN and MSegRNN with vary look-back window length
Our model exhibited smaller losses across various look-back windows, demonstrating the superiority of our method. Notably, the advantage of our method is more pronounced with smaller look-back windows. As the look-back window increases, this advantage diminishes, possibly due to noise introduced by segmentation affecting the model. Nevertheless, our model maintained a considerable advantage.
@@ 5.4 Ablation study
720 0.341 0.335 0.350 0.342 0.380 0.377 0.357 0.376
ETTh1 96 0.365 0.384 0.377 0.387 0.480 0.446 0.368 0.395
192 0.415 0.413 0.423 0.416 0.520 0.472 0.408 0.419
336 0.463 0.441 0.475 0.437 0.543 0.491 0.444 0.440
720 0.468 0.460 0.476 0.454 0.549 0.515 0.446 0.457
ETTh2 96 0.275 0.326 0.281 0.327 0.326 0.368 0.278 0.335
192 0.354 0.378 0.368 0.385 0.418 0.420 0.359 0.389
336 0.404 0.416 0.402 0.416 0.462 0.456 0.421 0.436
720 0.407 0.428 0.421 0.435 0.467 0.472 0.432 0.455
ETTm1 96 0.314 0.345 0.315 0.346 0.535 0.454 0.330 0.369
192 0.361 0.374 0.362 0.376 0.556 0.469 0.369 0.392
336 0.388 0.392 0.390 0.395 0.576 0.480 0.399 0.412
720 0.448 0.430 0.446 0.427 0.631 0.509 0.454 0.443
ETTm2 96 0.170 0.248 0.173 0.250 0.215 0.298 0.173 0.255
192 0.233 0.292 0.235 0.293 0.283 0.339 0.237 0.298
336 0.294 0.331 0.297 0.332 0.352 0.381 0.296 0.336
720 0.391 0.389 0.392 0.388 0.463 0.441 0.389 0.407
Average 0.335 0.355 0.341 0.357 0.425 0.411 0.340 0.372
@@ 5.5 Model efficiency
Since our Mamba structure has not undergone hardware-level optimization, this section focuses on comparing the model’s memory usage to verify its lightweight characteristic. The lightweight characteristic is primarily due to the use of a small number of parameters in our Mamba structure experiments. Using the Weather dataset as an example, the memory usage comparison with batch size 64 is shown in Table 4.
Table 4: Memory Usage of Two Methods
Method Memory Usage (GB)
96 192 336 720
MSegRNN 0.98 1.17 1.56 2.48
SegRNN 0.91 1.14 1.53 2.41
Compare +7.7% +2.6% +2.0% +2.9%
As shown in the table, the model’s exhibits only marginal increases in memory usage compared to SegRNN, affirming its lightweight characteristics.
@@ 6 Conclusion
Appendix
Appendix A Datasets
For our Long-term Time Series Forecasting (LTSF) study, we employed several widely-used multivariate datasets, including:
ETTs: This collection consists of electricity transformer data from two counties in China, spanning two years. The datasets include various subsets: ETTh1 and ETTh2 provide hourly data, and ETTm1 and ETTm2 offer 15-minute interval data. Each entry encompasses the target variable "oil temperature" alongside six power load indicators.
Electricity: This dataset contains hourly electricity consumption records for 321 clients from 2012 to 2014, offering insights into consumer energy usage patterns over time.
Weather: This dataset captures 21 meteorological variables, such as temperature and humidity, at 10-minute intervals throughout 2020. It serves as a comprehensive resource for weather prediction and climate studies.
These datasets are instrumental for benchmarking and advancing forecasting models in the context of long-term time series analysis.
Appendix B Code Framework
The implementation of our model framework is primarily derived from the official SegRNN code and the minimal implementation of Mamba. This includes data preprocessing and evaluation, model parameter settings and device selection, as well as result output. We sincerely thank the authors of these open-source frameworks and methods for their contributions to our experiments.
The official SegRNN code is available at:https://github.com/lss-1138/SegRNN.
The official minimal implementation of Mamba can be found at:https://github.com/johnma2006/mamba-minimal
Appendix C Detail Configuration
The main hyperparameter settings of our model are shown in Table 5. The parameter $d_{\text{state}}$ is a hyperparameter in the Mamba architecture, which is detailed in Mamba [Gu and Dao, 2023]. The parameter settings and results for other methods are taken from their original papers or official code repositories.
Table 5: Detail configuration of MSegRNN
Datasets d_state seg_len d_model learning_rate dropout batch_size
ETTh1 2 12 512 0.0003 0.1 256
ETTh2 4 24 512 0.0015 0.5 256
ETTm1 4 24 512 0.0010 0.1 256
ETTm2 4 24 512 0.0010 0.1 256
Weather 4 24 512 0.0007 0.1 64
Electricity 2 48 512 0.0012 0.01 128
We also employed a different learning rate adjustment strategy. The learning rate remained constant for the first 15 epochs and then gradually multiplied by 0.9. This approach actually increases the learning rate compared to the adjustment method used in SegRNN.
Appendix D Comparison with extra methods
We observed that some methods also utilized Mamba for information augmentation, we selected several long-term time series analysis models based on the Mamba architecture: Time-Machine[Ahamed and Cheng, 2024], Bi-Mamba4TS[Liang et al., 2024], and DT-Mamba[Wu et al., 2024], for comparison. The results are from official paper, respectivly . The result is shown in Table 6.
Table 6: Perfromance of different models
Datasets Models MSegRNN TimeMachine Bi-Mamba4TS DTMamba
MSE MAE MSE MAE MSE MAE MSE MAE
Weather 96 0.155 0.193 0.164 0.208 0.164 0.212 0.171 0.218
192 0.206 0.244 0.211 0.250 0.214 0.256 0.220 0.257
336 0.260 0.288 0.256 0.290 0.269 0.296 0.274 0.296
720 0.341 0.340 0.342 0.343 0.348 0.349 0.349 0.346
ETTh1 96 0.365 0.384 0.364 0.387 0.376 0.403 0.386 0.399
192 0.415 0.414 0.415 0.416 0.411 0.425 0.426 0.424
336 0.463 0.441 0.429 0.421 0.455 0.445 0.480 0.450
720 0.468 0.460 0.458 0.453 0.460 0.464 0.484 0.470
ETTh2 96 0.275 0.326 0.275 0.334 0.289 0.341 0.290 0.340
192 0.354 0.378 0.349 0.381 0.367 0.389 0.366 0.392
336 0.404 0.416 0.340 0.381 0.410 0.424 0.380 0.409
720 0.407 0.428 0.411 0.433 0.421 0.439 0.416 0.437
ETTm1 96 0.314 0.345 0.317 0.355 0.312 0.354 0.325 0.360
192 0.361 0.374 0.357 0.378 0.358 0.383 0.375 0.386
336 0.388 0.392 0.379 0.399 0.388 0.404 0.396 0.405
720 0.448 0.430 0.445 0.436 0.444 0.433 0.454 0.442
ETTm2 96 0.170 0.248 0.175 0.256 0.174 0.259 0.177 0.259
192 0.233 0.292 0.239 0.299 0.240 0.304 0.240 0.300
336 0.294 0.331 0.287 0.332 0.303 0.345 0.310 0.345
720 0.391 0.389 0.371 0.385 0.402 0.407 0.395 0.394
Electricity 96 0.149 0.242 0.142 0.236 0.142 0.238 0.166 0.256
192 0.165 0.259 0.158 0.250 0.157 0.253 0.178 0.268
336 0.181 0.275 0.172 0.268 0.172 0.271 0.197 0.289
720 0.218 0.307 0.207 0.298 0.200 0.297 0.243 0.326
It can be observed that our method performs notably well on the weather, ETTh2, and ETTm2 datasets, while the TimeMachine method shows significant advantages on the ETTh1 and Electricity datasets. Additionally, the Bi-Mamba4TS method also demonstrates a clear advantage on the Electricity dataset.